OpenStudy (summersnow8):

Use the fundamental identities to find the exact values of the remaining trig function of x: cos x= 1/ sqrt 10 and csc x= - sqrt (5) /2 the answers include: Tan x= -3 cot x= -1/3 sec x=sqrt 10 sin x= -3/ sqrt 10

OpenStudy (whpalmer4):

\[\cos x = \frac{1}{\sqrt{10}}\]Looking at the answer choices, notice that \[\sec x = \frac{1}{\cos x}\]If we take \[\sec x = \frac{1}{\cos x} = \frac{1}{\frac{1}{\sqrt{10}}} = 1*\frac{\sqrt{10}}{1} =\sqrt{10} \]

OpenStudy (summersnow8):

okay, i will c if i can figure out the others